OpenStudy (anonymous):

which point lies on the line defined by 3x+6y=2?

OpenStudy (anonymous):

to check just u pug the values of the point in place of x and y and if it satisfies the equation then it lies on the line otherwise it doesn't

OpenStudy (anonymous):

Just put it in y = mx + b form Subtract 3x over 6y = -3x + 2 Divide 6 y = -(1/2)x + 1/3 b is the y intercept (when x = 0). y = 1/3 When y = 0 Just use one of the above steps and plug in 0 for y, I'm going to use 6y = -3x + 2 0 = -3x+2 Simplify and solve for x 3x = 2 x = 2/3 So the points are (2/3, 0) and (0, 1/3) lie on the line.

OpenStudy (anonymous):

:) That's actually too much work. Just do what @matricked said. I'm going to leave my answer in case it's any use for you. Plug in random value for x and whatever y equals, that is a point on the line.